OpenStudy (anonymous):

how about this one what would the answer to be for the y intercept if the problem was 2y+3y+5=19+3x , i already have the slope which is 0

OpenStudy (saifoo.khan):

Whenever y intercpts, x = 0

OpenStudy (saifoo.khan):

Whenever x intercpts, y = 0

OpenStudy (angela210793):

5y=14+3x y=14/5+3x/5 slope=3/5